A cylindrical glass of water (H2O) has a radius of 5.66 cm and a height of 15.8 cm. The density of water is 1.00 g/cm3. How many moles of water are contained in the glass?
1
Expert's answer
2020-09-23T09:02:30-0400

The amount of substance (measured in moles) is

"\displaystyle \nu = \frac{m}{\mu}"

where "\mu (H_2O) = 2\cdot A_r(H)+ A_r(O) = 2 \cdot 1 +16 = 18" g/mol.

Mass can be found from

"m = \rho \cdot V = \rho \cdot \pi r^2 h = 1 \cdot 3.1415 \cdot 5.66^2 \cdot15.8 = 1590" g

So,

"\nu = \frac{1590 \;g}{18 \; g/mol}= 88.3 \; mol."

Answer: 88.3 mol